OpenStudy (anonymous):

Find the constant of variation k for the direct variation. 2x + 6 y =0

ganeshie8 (ganeshie8):

k is just the slope of the given equation. rearrange the equation in the form : \(\huge y = mx + c\) then m is the constant of variation. can you rearrange the equation in above form.. ?

OpenStudy (anonymous):

find the constant of variation k for the direct variation. 4x = -6y

OpenStudy (unklerhaukus):

k=m

OpenStudy (anonymous):

y=2/3x+0 so constant variation of k = 2/3